Assume that Bay King Products sells three varieties of canned seafood with the following prices costs: Selling Price per Case $125 $150 $100 Variable Cost per Case $95 100 Fixed Cost per Month Premium Royal Deluxe Entire firm 80 $246,500 The sales mix (in cases) is 30% Premium, 20% Royal, and 50% Deluxe. Required What is the weighted average contribution margin? 2. What is the break even point in total monthly sales units? 3.What is the breakeven point in sales revenues? 4. How many total units do they need to sell to earn $20,000 per month? 5. If the tax rate is 20% how many total units do they need to sell to earn $20,000 after-tax? 6.Total sales average 9,000 cases per month and range seasonally from 8,000 to 10,000 cases. What is the expected annual profit before taxes? Bay Kings General Manager believes an advertising campaign costing $150,000 annually would increase monthly sales by 600 total cases per month. What would be the annual profits before taxes if Bay King undertook this investment? 7. 8. Due to a change in market conditions the variable cost of a case of the Royal product is expected to 9.Base your answer on the original data for Bay King. If labor cost increases cause the cost of all products 10. In order to address the decline in profits resulting from the increased labor costs described in question 9 increase by 15%. What will the weighted average contribution margin equal if this change occurs? to increase by two dollars per case what is the new expected level of monthly profits before taxes? Bay King is considering increasing the price of all products by three dollars per case. They expect the impact of this across the board price increase to be a decrease in the average number of total cases sold per month. Given the changes to costs and prices, how many cases per month must they sell to earrn profits equal to their original monthly profits based on 9,000 total cases?Solve question 6 to 10
Answer #1

Ans.     Selling price             Variable cost          Contribution   P/V ratio     Sales ratio Weighted contribution margin

Premium            125                          95                           30                 24%            30%                 7.2%

Royal                 150                        100                           50                33.33%        20%                 6.67%

Deluxe               100                         80                            20                20%             50%                 10%

Total    375    275    100           26.67%                                 23.87%

Weighted average contribution margin = 23.87%

2. Break even sales in units = 246500/100 = 2465 unitts

3. Break ever points sales in value = 246500/.2387 = $1032677

4. Total no of units need to sale if required profit $20000 = (246500+20000)/100 = 2665 units

5. if tax rate is 20% then no of units required to sell for earn $20000 = 20000/.80 =(25000+246500)/100=2715 units

6. Contribution if no of units 9000 (9000X100)*12 = 10800000

Fixed cost                                                              = 246500

Profit before tax                                                     = 10553500

7. due to advertisement no of sales unit increase is 600 per month

Total additional contribution (600X100)X12 = 720000

Less: additional advertisement exp             = 246500     

Profit before tax                                          = 473500

8. Calculation of revised weighted average contribution margin

Total sales value                 = 375

Less:variable cost

Premium   =                           (95)

Royal        =100X1.15 =         (115)

Deluxe      =                            (80)

Contribution margin            = 85   

Contribution margin = 85/375X100 = 22.67%

9. in the case of bay king if labour cost of all product is increase by 2 dollar , calculation of profit before tax

Revised variable cost

Premum    = 95+2   = 97

Royal       = 100+2 = 102

Deluxe      = 80+2 = 82

Total variable cost = 281

Sale value = 375

Revised contribution (375-281) = $94

Total contribution per month (600X94) = 56400

Less: Fixed cost per month

         (246500+150000)/12                   = 33042

Profit before tax before tax    = 23358

10. Revised sale price = 375+3 = 378

Less variable cost                 = 281

Contribution per unit =   97

Total contribution for 9000 cases (97X9000) = 873000

Less: fixed cost (246500+150000) = 396500

profit before tax    = 476500

profit before tax per month = $39708